# Just-In-Time Objectives: A General Approach for Specialized AI Interactions

### v1

@@ 5.2.2. Are Poppins outputs useful?
Participants find system outputs useful for their tasks, particularly with the Poppins-experts system, which was deemed “useful” or “very useful” by 82.4% of participants (Baseline: 64.7%, Poppins-tools: 64.7%) and “good quality” or “very good quality” by 88.2% of participants (Baseline: 64.7%, Poppins-tools: 58.8%).
@@
The induced JIT objectives effectively steered downstream generation, producing outputs that were more useful than those yielded by the baseline condition. Users attributed their preference of Poppins output to the presence of experts, which produced more specific, opinionated, and ultimately useful outputs. For example, P11, a PhD student in chemical engineering, found Poppins outputs comparable to Deep Research outputs using only a fraction of a time, and they felt Poppins results were more reliable and accurate. Similarly, P19, a bioengineering PhD student working on a research proposal, was impressed at the quality of the Poppins output that they felt was “much more in depth, even more than o1” (one of OpenAI’s advanced reasoning models):
“[Poppins] definitely gave more in-depth analysis of my input, and also the answers it gave me were very detailed—I would say PhD-level answers […] Some of the suggestions are actually what my advisor or reading committee, who are [redacted university] professors, told me before. So it was very surprising.” — P19
These outputs often broke participants’ expectations of what language models could produce. When reading feedback on their short story from a Character Consistency Specialist, P30 was pleasantly surprised by the expert’s suggestion to introduce ambiguity and inconsistency into the protagonist’s character arc to increase story complexity: “I don’t think an AI would ever tell me to make my characters inconsistent, but this expert did!” Others found that Poppins outputs nicely augmented their abilities on tasks they find challenging. When translating manga, P16 expressed that the expert feedback helped them incorporate more creative liberty while preserving semantic accuracy in their translation: “I’m not the most creative writer, so I think having these ideas stretch what I’m capable of doing.”
@@
Importantly, participants found that system-generated experts also supported better interpretation and critical evaluation of system outputs. Conventional AI systems blend a myriad of sources and perspectives into a single generic voice, obscuring users from contextual information and making it difficult to track the validity of outputs:
“I really like the different [expert] perspectives. With normal AI, they just combine it all—you don’t know who’s saying what. Then it gets rid of our ability to get context on who’s writing this opinion… and their biases. AI just tries to generalize it into this all-in-one writing guide.” — P30
@@ 5.3. Study Limitations
For consistency across participants on an in-depth task, this study evaluates on writing tasks, so findings may not generalize to other tasks. Participants only applied the tool for one task; while we randomized the order of baseline and Poppins conditions to control for order effects, it may be valuable to evaluate on independent tasks and in organic, longitudinal task settings. Our study sessions found that users are more familiar with chat interfaces like ChatGPT and have grown accustomed to prompting, even with its inconveniences. Further evaluations may explore the cognitive load Poppins incurs by having users engage with objectives and intermediate tool design decisions, which may inform design guidance on the appropriate balance between user effort and resulting output quality with JIT objectives.
@@ 6.1.1. Expanding the scope of just-in-time objectives
Future work could also explore using JIT objectives to not only augment prompts, but perform model finetuning for more robust behavior or train a cheaper distilled model for reuse (Hu et al., 2022; Wu et al., 2024a; Xu et al., 2024). Within the space of prompt interventions, future work could explore test-time scaling approaches with enhanced verifiers to ensure model alignment with stated objectives (Snell et al., 2024; Beirami et al., 2024). The current JIT objective applies the human-readable specification in its prompt augmentations, but this need not be the case: we might achieve higher performance by optimizing over alternative prompt formulations (Khattab et al., 2024) to best achieve the same stated objective.
Our larger-scale experiment provided preliminary evidence that just-in-time objectives could work effectively in a number of domains beyond those we had anticipated. However, we would benefit from a deeper understanding of the performance characteristics of our method not just in general, but for specific task domains. What are domains where it is especially prone to failure? Future work can investigate this question in more depth. Lastly, just-in-time objectives are formulated as objectives for an individual user in our work, but many important objectives extend beyond an individual user. Novel techniques might observe the joint activity of groups of individuals, infer shared objectives, and apply these objectives to appropriately steer AI assistance within a multi-user context.
@@ 6.1.2. JIT objectives and human-AI interaction
Our work on generative UIs also intersects with ongoing debates about the role of HCI and design as AI advances further. Dominant AI narratives tend to automate away the work of HCI because they envision a model that absorbs and obscures its inner workings, including interface design decisions. By contrast, our just-in-time objectives architecture makes visible the numerous design decisions that go between an AI system and a user interface. The architecture acknowledges that there is a vast design space of potential model capabilities to surface, and lacking a singular “right answer,” HCI and design expertise is critical to navigate these decisions. If successful, just-in-time objectives might shift the borders of this relationship. Designers would still need to understand user needs, but they might elicit hyperlocal needs or use custom objective induction approaches. They would still design and iterate on interactive systems, but they might build them at a meta level by authoring custom generators, evaluators, and more complex architectures built of these elements. These design choices still fundamentally depend on having domain expertise, a rich understanding of users, and the ability to design the right components (i.e., the areas in which HCI excels). Thus, we envision that HCI work will continue to have a strong impact even as AI advances, and we advocate for more designer involvement here, not less.
@@ 6.3.2. Privacy
JIT objectives require observing the user’s current state. In our evaluation, this observation was always done with explicit participant consent. If JIT objectives were to be spread further, however, they might encounter violations of expected privacy norms (Nissenbaum, 2004). Users are more likely to expect that an application can see what the user is doing within it, but a broader view at the browser or operating system level could be less expected. As a result, users may accidentally divulge information to the application that they wish it did not see. Tools will need to exist for users to pause recording, clear history, and otherwise build contextual integrity modules (Shaikh et al., 2025b) that are careful about what they see and remember.
@@ Acknowledgements.
We are grateful to Yijia Shao for her contributions to early versions of this project that shaped our current ideas. We also thank Dora Zhao, Farnaz Jahanbakhsh, and Poonam Sahoo for their insightful feedback on the paper. We thank our study participants for sharing their valuable perspectives on our research.
@@ A.6. UI code generation
- Please use the attached museService.js file to help you generate the tool. Use the functions to retrieve the entities and make LLM calls rather than using hard-coded data to populate the interface.
- Use this import: import {{ museService }} from \$lib/museService’;‘
@@ A.7. UI code critique
Appendix B User Evaluations
We include survey questions, interview questions, and participant background information for our user evaluations.
B.1. Surveys
B.1.1. Study 1 Survey
@@ B.1.2. Study 2 Survey
B.1.3. Poppins Study Survey
@@ B.2. Interview
  • What was your general impression of the output generated in both conditions?
  • What was your impression of the goals generated in [Poppins task]?
  • What was your impression of the selected tool format and tool expertise generated in [Poppins task]?
  • How did you find the process of trying to improve the system-generated assistance in both conditions?
  • What was your overall impression of the systems you used in each condition? Anything that excited you or surprised you?
@@
  • What was your general impression of the tool that the system generated?
  • What was your impression of the goals?
  • What was your impression of the selected tool format and tool expertise?
  • What was your overall impression of this design direction? Anything that excited you or surprised you?
  • Are there other tasks where you would like to try out our tool? Why? Would you want to use a tool like this in your everyday life?
B.3. Participant Background

### v2

@@ 5.2.2. Are Poppins outputs useful?
Participants find system outputs useful for their tasks, particularly with the Poppins-experts system, which was deemed “useful” or “very useful” by 82.4% of participants (Baseline: 64.7%, Poppins-tools: 64.7%) and “good quality” or “very good quality” by 88.2% of participants (Baseline: 64.7%, Poppins-tools: 58.8%) (Figure 13). Directly comparing individual participant ratings for Poppins-experts and Baseline on a -3 to +3 rating scale (Figure 14), we find that Poppins-experts achieves generally higher ratings for overall quality (+0.59; Baseline: 1.41, Poppins-experts: 2.0), user control (+0.35; B: 1.18, P: 1.53), usefulness (+0.23; B: 1.65, P: 1.88), and relevance (+0.11; B: 2.18, P: 2.29). Though this study was designed for qualitative analyses, we conduct an exploratory statistical analysis. A Wilcoxon signed-rank test indicates that for overall quality, participant ratings are significantly lower for the Baseline than for Poppins-experts ($W=8.0$, z=1.75z=-1.75, $p<.05$); we find no significant differences for user control, usefulness, and relevance at the current sample size.
@@
The induced JIT objectives effectively steered downstream generation, producing outputs that were more useful than those yielded by the baseline condition. Users attributed their preference of Poppins output to the presence of experts, which produced more specific, opinionated, and ultimately useful outputs. For example, P11, a PhD student in chemical engineering, found Poppins outputs comparable to Deep Research [^4] Deep Research is an OpenAI agent that conducts multi-step research on the internet. outputs using only a fraction of a time, and they felt Poppins results were more reliable and accurate. Similarly, P19, a bioengineering PhD student working on a research proposal, was impressed at the quality of the Poppins output that they felt was “much more in depth, even more than o1.” [^5] OpenAI o1 is an advanced reasoning model released in late 2024.
“[Poppins] definitely gave more in-depth analysis of my input, and also the answers it gave me were very detailed - I would say PhD-level answers […] Some of the suggestions are actually what my advisor or reading committee, who are Stanford professors, told me before. So it was very surprising.” — P19
These outputs often broke participants’ expectations of what language models could produce. When reading feedback on their short story from a Character Consistency Specialist, P30 was pleasantly surprised by the expert’s suggestion to introduce ambiguity and inconsistency into the protagonist’s character arc to increase story complexity: “I don’t think an AI would ever tell me to make my characters inconsistent, but this expert did!” Others found that Poppins outputs augmented their abilities on tasks they find challenging. When translating manga, P16 expressed that the expert feedback helped them incorporate more creative liberty while preserving semantic accuracy in their translation: “I’m not the most creative writer, so I think having these ideas stretch what I’m capable of doing.”
@@
Participants found that system-generated experts supported interpretation and critical evaluation of system outputs. Conventional AI systems blend a myriad of sources and perspectives into a single generic voice, obscuring users from contextual information and making it difficult to track the validity of outputs:
“I really like the different [expert] perspectives. With normal AI, they just combine it all - you don’t know who’s saying what. Then it gets rid of our ability to get context on who’s writing this opinion… and their biases. AI just tries to generalize it into this all-in-one writing guide.” — P30
@@ 5.3. Study Limitations
For consistency across participants on an in-depth task, this study evaluates on writing tasks, so findings may not generalize to other tasks. Participants only applied the tool for one task; while we randomized the order of baseline and Poppins conditions to control for order effects, it may be valuable to evaluate on independent tasks and in organic, longitudinal task settings. Our study sessions found that users are familiar with chat interfaces like ChatGPT and have grown accustomed to prompting, even with its inconveniences. Further evaluations may explore the cognitive load Poppins incurs by having users engage with objectives and intermediate tool design decisions, which may inform design guidance on the appropriate balance between user effort and output quality with JIT objectives.
@@ 6.1.1. Expanding the scope of just-in-time objectives
Future work could also explore using JIT objectives to not only augment prompts, but perform model finetuning for more robust behavior or train a cheaper distilled model for reuse (Hu et al., 2022; Wu et al., 2024c; Xu et al., 2024). Within the space of prompt interventions, future work could explore test-time scaling approaches with enhanced verifiers to ensure model alignment with stated objectives (Snell et al., 2024; Beirami et al., 2024). The current JIT objective applies the human-readable specification in its prompt augmentations, but this need not be the case: we might achieve higher performance by optimizing over alternative prompt formulations (Khattab et al., 2024) to best achieve the same stated objective.
Our larger-scale experiment provided preliminary evidence that just-in-time objectives could work effectively in a number of domains beyond those we had anticipated. However, we would benefit from a deeper understanding of the performance characteristics of our method not just in general, but for specific task domains. What are domains where it is especially prone to failure? Future work can investigate this question in more depth. Lastly, JIT objectives are formulated as objectives for an individual user in our work, but many important objectives extend beyond an individual user. Novel techniques might observe the joint activity of groups of individuals, infer individual and shared objectives, and apply these objectives to appropriately steer AI assistance within a multi-user context.
@@ 6.1.2. JIT objectives and human-AI interaction
Our work on generative UIs also intersects with ongoing debates about the role of HCI and design as AI advances further. Dominant AI narratives tend to automate away the work of HCI because they envision a model that absorbs and obscures its inner workings, including interface design decisions. By contrast, our just-in-time objectives architecture makes visible the numerous design decisions that go between an AI system and a user interface. The architecture acknowledges a vast design space of potential model capabilities to surface, and lacking a singular “right answer,” HCI and design expertise is critical to navigate these decisions. If successful, JIT objectives might shift the borders of this relationship. Designers would still need to understand user needs, but they might elicit hyperlocal needs or use custom objective induction approaches. They would still design and iterate on interactive systems, but they might build them at a meta level by authoring custom generators, evaluators, and more complex architectures built of these elements. These design choices still fundamentally depend on having domain expertise, a rich understanding of users, and the ability to design the right components (i.e., the areas in which HCI excels). Thus, we envision that HCI work will continue to have a strong impact as AI advances, and we advocate for more designer involvement, not less.
@@ 6.3.2. Privacy
JIT objectives require observing the user’s current state. In our evaluation, this observation was always done with explicit participant consent. If JIT objectives were to be spread further, however, they might encounter violations of expected privacy norms (Nissenbaum, 2004). Users are more likely to expect that an application can see what the user is doing within it, but a broader view at the browser or operating system level could be less expected. As a result, users may accidentally divulge information to the application that they wish it did not see. Tools will need to exist for users to pause recording, clear history, and build contextual integrity modules (Shaikh et al., 2025b) that are careful about what they see and remember.
@@ Acknowledgements.
We are grateful to Yijia Shao for her contributions to early versions of this project that shaped our current ideas. We also thank Dora Zhao, Farnaz Jahanbakhsh, Helena Vasconcelos, and Poonam Sahoo for their insightful feedback on the paper. We thank our study participants for sharing their valuable perspectives on our research. This work was supported by the Stanford Institute for Human-Centered Artificial Intelligence (HAI) and NSF Award 2402647. Michelle Lam was supported by a Stanford Interdisciplinary Graduate Fellowship.
@@ A.6. UI code generation
- Please use the attached service.js file to help you generate the tool. Use the functions to retrieve the entities and make LLM calls rather than using hard-coded data to populate the interface.
- Use this import: import {{ service }} from \$lib/service’;‘
@@ A.7. UI code critique
Appendix B Poppins Tools
We include additional samples of tools generated by Poppins from study participants (Figure 15) and the research team (Figure 16).
Figure 15. Additional examples of Poppins-tools output on participant writing tasks demonstrate various forms of interactive assistance. A Concept Visualizer creates diagrams for theoretical concepts, a Cultural Perspective Highlighter aids iteration on a scholarship essay, an Academic Jargon Organizer surfaces inline wording improvements on a presentation script, an Academic Paper Finalizer creates an interactive checklist of paper to-dos, and a Hypothesis-Evidence Mapper links hypotheses and evidence in research notes.
Additional examples of Poppins-tools generations for participant writing tasks. Five screenshots show generated tools. A Concept Visualizer creates visual diagrams for psychology concepts. A Cultural Perspective Highlighter highlights cross-cultural themes within an essay. An Academic Jargon Organizer proposes inline suggestions to improve a presentation script. An Academic Paper Finalizer has a checklist of paper to-dos and a summary progress bar that updates. A Hypothesis Evidence Mapper displays extracted hypotheses from research notes and interactively links them to evidence and planned experiments.
Figure 16. Poppins can produce tailored tools to support visual design tasks (Scientific Color Palette Generator), multi-user coordination (Camping Destination Dashboard), and performing arts contexts (Character Arc Visualizer). The system can generate tools like a customized text editor where suggestions can be proposed and applied directly inline (Transition Highlighter).
Additional Poppins-tools demonstrations. Four screenshots show generated tools. A Scientific Color Palette Generator that displays palette options for different selected data types. A Camping Destination Dashboard shows camping sites with photos, a summary of research progress for each, and functionality to check availability and calculate total cost. A Character Arc Visualizer allows the user to select a character from the script and view an Emotional & Psychological analysis, Dialogue & Speech Patterns, and Character Development Timeline. A Transition Highlighter produces a text editor that highlights transitions and provides suggestions to improve them inline.
Appendix C User Evaluations
We include survey questions, interview questions, and participant background information from our user evaluations.
C.1. Surveys
C.1.1. Study 1 Survey
@@ C.1.2. Study 2 Survey
C.1.3. Poppins Study Survey
@@ C.2. Interview
  • Output. What was your general impression of the output generated in both conditions?
  • Goals. What was your impression of the goals generated in [Poppins task]?
  • Components. What was your impression of the selected tool format and tool expertise generated in [Poppins task]?
  • Editing process. How did you find the process of trying to improve the system-generated assistance in both conditions?
  • Overall comparison. What was your overall impression of the systems you used in each condition? Anything that excited you or surprised you?
@@
  • Output. What was your general impression of the tool that the system generated?
  • Goals. What was your impression of the goals?
  • Components. What was your impression of the selected tool format and tool expertise?
  • Overall comparison. What was your overall impression of this design direction? Anything that excited you or surprised you?
  • Are there other tasks where you would like to try out our tool? Why?
  • Would you want to use a tool like this in your everyday life?
  • Did you have any other thoughts or reflections from the study? Anything that surprised you? Feedback to share?
C.3. Participant Background